What is the slope of x^2-9=0, x^2+5=0, 4x^2=0, 2x^2-8=0, x^2+16=0, x^2+7=7,
Answer by Alan3354(69443)

What is the slope of x^2-9=0, x^2+5=0, 4x^2=0, 2x^2-8=0, x^2+16=0, x^2+7=7
----------------
What is the slope of x^2-9=0
The slope at any point is x*2a
m = 2x
------
4x^2=0
m = 8x
-------
Do the others in the same way.